Required information

[The following information applies to the questions displayed below.]

The following data pertain to the Oneida Restaurant Supply Company for the year just ended.

Budgeted sales revenue $ 205,000
Actual manufacturing overhead 338,000
Budgeted machine hours (based on practical capacity) 8,000
Budgeted direct-labor hours (based on practical capacity) 20,000
Budgeted direct-labor rate $ 13
Budgeted manufacturing overhead $ 364,000
Actual machine hours 11,000
Actual direct-labor hours 18,000
Actual direct-labor rate $ 17

Required:

1. Compute the firm’s predetermined overhead rate for the year using each of the following common cost drivers: (a) machine hours, (b) direct-labor hours, and (c) direct-labor dollars. (Round your answers to 2 decimal place.)
Answer #1

a)

Predetermined overhead rate using machine hours = Budgeted manufacturing overheads/Budgeted machine hours

= 364,000/8,000

= $45.5 per machine hour

b)

Predetermined overhead rate using direct labor hours = Budgeted manufacturing overheads/Budgeted direct labor hours

= 364,000/20,000

= $18.2 per direct labor hour

c)

Budgeted direct labor dollars = Budgeted direct labor hours x Budgeted direct-labor rate

= 20,000 x 13

= $260,000

Predetermined overhead rate using direct labor dollars = Budgeted manufacturing overheads/Budgeted direct labor dollars

= 364,000/260,000

= $1.4 per direct labor dollar

Kindly comment if you need further assistance. Thanks